# The representation of Convectively Coupled Equatorial Waves and upscale energy transfer in models with explicit and parametrized convection

### v1

@@ 4.1 Simulated waves and precipitation
From this first analysis of the waves and precipitation in the RAL and GAL models, we have observed that both models simulate more rain than the observations, the RAL particularly so. Excessive tropical rain is commonly observed in CPMs (e.g. Howard et al. 2024). The GAL model doesn’t capture the high variability of precipitation seen in the tropics, whereas the RAL does. Again, this has been previously observed by other CPM studies (e.g. Judt and Rios-Berrios 2021). Regarding equatorial wave propagation, the waves appear comparable on a Hovmöller plot but a closer investigation reveals there are differences in simulated propagation speed between the analysis and models, with both models simulating a too fast Kelvin and WMRG wave, and GAL with a higher variability of simulated wave speeds. The Rossby wave speeds in GAL are comparable to the analysis, and faster in RAL with higher variability. The distribution of wave variability in the study period across the equatorial band and full depth of the troposphere is comparable between all models and the analysis.
@@ 4.2 Simulated upscale energy transfer and dynamics
To get a first indication of the reasons for the differences in simulated propagating convection between the models, time mean maps of interscale energy transfer for $L$ = 440 km at 200 hPa (the top of the tropical troposphere) are shown in Fig. 4 for GAL and RAL. This is the level where horizontal divergence is at a maximum, and so we expect to see high values of upscale energy transfer, because of the mechanism explained at the end of this subsection. For the interscale energy transfer analysed in this study, we do not compare the models to the analysis product. This is because it is at a much coarser resolution than the models and is driven by a data assimilation (DA) scheme with parametrized convection. This means we would expect to see low magnitudes of interscale energy transfer in the anlaysis that we would not trust at the lower length scales analysed since this would be too close to the grid scale of the analysis. Furthermore, the presence of DA in the analysis product would not allow for the investigation of propagating features developing on their own for a longer time period without the impact of the DA scheme.
In Fig. 4, the blue colours indicate the energy is being transferred upscale from scales below 440 km to scales larger than this. The red colours indicate that energy is being transferred downscale from scales larger than 440 km to scales smaller than this. One clear and obvious difference between the two is that in RAL, there is orders of magnitude more upscale energy transfer, and this is reflected by the deep-tropical mean at 200 hPa being net upscale. The mean was taken over 15S to 15N to avoid impact of sub-tropical systems, such as the subtropical jet. The domain 20S-20N is shown to allow for visualisation of the full extent of the ITCZ in Boreal Summer. The values shown are broadly consistent with those computed by Faranda et al. (2018), the only other paper that has studied local interscale energy transfers in the atmosphere. The magnitudes reported here are on the same order as atmospheric turbulent dissipation rates calculated from observations (Nowak et al., 2025).
@@
The ITCZ is visible in Fig. 4 as the blue band of upscale centred around 10N. Similarly, the SPCZ is visible extending west from New Guinea to the other side of the dateline. Both convergence zones are associated with much stronger upscale energy transfer in the RAL model. Another key feature in Fig. 4 is the abundant upscale energy transfer in the Indian Ocean. This period coincides with a strong negative IOD event (cf. Table 1), which is associated with enhanced convection in the southeast Indian Ocean (Saji et al., 1999) – a region which is convectively active even without the negative IOD. Once more, this region of upscale energy transfer is orders of magnitude greater in RAL than GAL, saturating the colour scale. Another upscale feature which is much larger and more extensive in RAL is the deep convection over the Amazon. Panel (c) in Fig. 4 provides further evidence that the upscale energy transfer at 200hPa is located at the tropical convergence zones. These zones are characterised by consistently high precipitation, and are reflected in the map of mean rainfall over the study period in this panel. The precipitation fields are from GPM-IMERG, and we can see from Fig. 4 that the RAL model is simulating a strong ITCZ in terms of 200hPa upscale energy transfer.
@@
Fig. 5 shows longitudinally-averaged, time-mean interscale energy transfer (shading, zero contour thick grey line), zonal mean zonal winds (black contour lines), and the mean meridional circulation (vectors). The main features of Fig. 5 which connect back to Fig. 4 are the sections of upscale energy transfer at the top of the troposphere centred around 200 hPa. This area of upscale is associated with time-average flow and divergence in the tropical upper troposphere as a result of deep convection and associated vertical velocities in the ITCZ centred around 10N. Near the surface, downscale transfer is seen associated with convergence as part of the ITCZ circulation. The RAL exhibits both stronger upscale and downscale energy transfer associated with the ascending branch of this circulation, suggesting that explicit representation of convection leads to RAL simulating interscale energy transfer that is significantly larger than that in the GAL model. The large scale Hadley circulation is of similar strength in both RAL and GAL, but RAL shows an enhanced secondary circulation at 500 hPa associated with convection. Centred around 5S and 150 hPa is a region of upscale energy transfer which can be identified with the region of strong upscale energy transfer over the Indian Ocean seen in Fig. 4. This region of strong upscale is also associated with downscale energy transfer at low levels underneath the deep convection associated with net flow convergence. As also seen in Fig. 4, this region is about twice as large in RAL compared to GAL. This is further evidence of the CPM’s ability to simulate strong upscale energy transfer associated with deep convection. At the southern edge of (a) in the upper troposphere, the signal of the subtropical jet is seen associated with strong zonal winds and strong downscale energy transfer.
@@
In the middle troposphere, there are two areas of weak upscale energy transfer in the RAL, but only one in the GAL. After consulting maps of time-mean interscale energy transfer at 500 and 600 hPa (not shown), the feature at this level in the RAL around 5-10N is associated with the stronger secondary ITCZ circulation in this model, and the presence of cumulus congestus associated with the West African Monsoon. This upscale feature is not present in the GAL. At around 700 hPa between 5 and 10S is another upscale feature which appears in both models. On consultation of the maps at this level this appears to be collocated with upscale energy transfer over the Amazon and Congo rainforests, possibly associated with shallow convection in these regions.
@@
Moving away from the time-mean picture, a Hovmöller of interscale energy transfer is shown in Fig. 6. The contour limits are 5 times higher for the 200 hPa panels than the 850 hPa panels due to the larger overall amount of kinetic energy due to higher wind speeds at this altitude. For the region shown in Fig. 6 spanning the Indian Ocean, Maritime Continent, and West Pacific, there are abundant eastward-propagating downscale energy transfer features at 850 hPa which are broadly collocated with the propagating rainfall and Kelvin waves in Fig. 2. These downscale features are occasionally coincident with upscale energy transfer at 200 hPa (e.g. see in the RAL starting at 90E on 2016-08-01), indicative of propagating deep convection (likely a CCKW). The picture at 200 hPa is mostly dominated by westward-propagating features. These Hovmöller diagrams were analysed with waves overlain and for symmetric, anti-symmetric, and off-equatorial interscale energy transfer (not shown). We found that the westward-propagating features were indeed indicative of CCEW activity: R1 waves in the case of the off-equatorial interscale energy transfer (0-20N average), and WMRG waves for the anti-symmetric part (calculated by taking the average of the difference between the fields in the 20 degrees north and south of the equator).
Figure 6: Time-longitude (Hovmöller) diagrams of $\mathcal{D}_{L}(\mathbf{u})$ for $L$ = 440 km averaged between 10S and 10N for the 20 day study period, limited to 45 - 180E to remove the strong orographic effects of the East African Highlands at 850 hPa. (a) for GAL at 200hPa; (b) RAL at 200 hPa; (c) GAL at 850 hPa, and (d) RAL at 850 hPa. Note the contour limits are 5 times higher in (a) and (b) than in (c) and (d). Colour-coded numbers in the top right of each panel are mean values of the data shown with the colour indicating net upscale (blue) or net downscale (red). The dashed line indicates the time of the snapshot profile shown in Fig. 7. The time mean has been removed to clarify the propagating features, and this has been done for Fig. 7, and for calculating the composites shown in Fig. 8.
@@ 4.3 Equatorial waves and upscale energy transfer
The picture of the imprint of the general circulation on the interscale energy transfer can be seen in Fig. 6 by the fact that the 200 hPa regions are net upscale and the 850 hPa regions net downscale. Once again it is evident that RAL is simulating significantly more upscale than GAL at 200 hPa – 2 orders of magnitude more in this case. To investigate how propagating features in Fig. 6 are associated with different equatorial waves, longitude-pressure snapshots at the time indicated by the dashed line in Fig. 6 of dynamical wave fields and interscale energy transfer are shown in Fig. 7. All wave snapshots are shown across the longitudes where they are strongest in this season. For the KW this is over the Indian Ocean, for the R1 wave it is over the Atlantic, and for the WMRG over the West Pacific.
From figures 7 (a) and (d), we can see upscale energy transfer related to Kelvin wave divergence. Upscale energy transfer in the upper troposphere in the divergent phase of the wave is observed because the convecting plume coupled to the wave is expanding near the tropopause, transferring energy from scales smaller than 440 km (the scale of the updraft) to scales larger than 440 km (toward the scale of the wave). This is balanced by downscale energy transfer in the convergent phase of the wave, both beneath the divergence in the lower troposphere, and laterally in the wave troughs either side of the upper-level wave divergence. Another thing to note from Fig. 7 is that the vertical structure of the Kelvin wave in RAL is closer to observations and theory than GAL (cf. Straub and Kiladis 2003; Yang et al. 2007a). The wave is higher amplitude, exhibits clear westward tilt with height, and upward propagation into the stratosphere in RAL, whereas the GAL does not exhibit these qualities as clearly, especially in the lower troposphere.
@@
The composites in Fig. 8 provide more information than the instantaneous structures seen in Fig. 7, demonstrating the systematic features that are present over time in interscale energy transfer and dynamical equatorial wave fields. Regarding the Kelvin wave and symmetric eastward-propagating interscale energy transfer, it can be seen that the dynamical wave composite is similar in both the RAL and the GAL models. As expected, there is more extreme magnitudes of upscale energy transfer in the RAL, and this appears to be more highly correlated with the KW divergence in the upper troposphere than the GAL. The correlations between the propagating interscale energy transfer and corresponding equatorial waves for varying $L$ are outlined in Fig. 9.
@@
Figures 8 (b) and (e) show again the positive correlation between R1 vorticity and interscale energy transfer. This positive correlation between off-equatorial interscale energy transfer and R1 vorticity is reflected in Fig. 9, with correlations in RAL being on average 10% higher than in GAL, and more consistent (same sign for all $L$).
@@
The WMRG composites span a greater longitudinal extent than the other waves in Fig. 8 (c) and (f). This is to allow room to show the striking differences in their simulation between the models. In the RAL there is the characteristic vertical propagation of the WMRG wave typically seen during the westerly phase of the QBO (Yang et al., 2012). The QBO is indeed in a westerly phase during DYAMOND Summer, see Table 1. No such vertical propagation is seen in the GAL. The GAL does however have a slightly stronger baroclinic wave structure, even if the wave is much weaker. The composite is again consistent with the snapshot for RAL, with southward WMRG winds being collocated with anti-symmetric upscale energy transfer. This collocation is also present in GAL albeit with lower values of interscale energy transfer.
@@
Covariance and correlations between symmetric eastward, off-equatorial westward, and anti-symmetric westward interscale energy transfer and the dynamical waves are presented for varying length scales in Fig. 9. The correlations and covariances displayed here are comparing the dynamical wave and interscale energy transfer composites shown in Fig. 8 for multiple length scales. For reference, the fifth data points for each line represent the $L=440$ km length scale plotted in Fig. 8. The composites for length scales 30 km up to 440 km were calculated on the N120 grid, with larger $L$ being calculated on the half degree data to save computational expense.
@@
Figure 9: Covariances and correlations between the fields shown in Fig. 8 as a function of characteristic length scale $L$. Kelvin wave divergence about the equator is compared with equatorially symmetric KW-following interscale energy transfer, R1 off-equatorial vorticity with off-equatorial R1-following interscale energy transfer, and WMRG cross-equatorial $\upsilon$ with anti-symmetric WMRG-following interscale energy transfer, as in Fig. 8. Correlations and covariances for GAL in cyan, for RAL in orange. The blue vertical lines represent multiples of the composite dominant wavelength, with the dotted line $\lambda/8$, dot-dash $\lambda/4$, dashed $\lambda/2$, and solid line at the dominant wavelength, $\lambda$.
@@
For the Kelvin wave in Fig. 9 (a) and (d), we can see that the covariance between the dynamical wave and interscale energy transfer is stronger in the RAL for length scales 880 km and lower (down to 30 km in these diagrams), comparable and at a maximum for $L=1760$ km, and then the other side of the zero line for the maximum length scale analysed here of 3410 km. Consultation of the composites at these two large length scales (not shown) reveals that the reason for this is because at $L=1760$ km, the regions of strong interscale energy transfer are close to the scale of the wave (between $\lambda/4$ and $\lambda/2$) and almost in-phase with them. This means a relatively high correlation of around 0.6 in both models, and with a highly negative covariance - negative (upscale) collocated with positive (divergent) wave-dynamical flow. However at the 3410 km length scale, the interscale energy transfer features are double the size, and are almost large enough to stretch across an entire wavelength with the same sign, leading to correlations close to 0 and a positive covariance, since the deepest regions of upscale energy transfer are located in the convergent phase of the wave. Correlations are strongest for length scales between 440 and 1760 km. The reasons for this will be discussed at the end of this section in conjunction with the information from the other waves. Correlations are slightly larger in magnitude in the RAL for 440 and 880km, and comparable for both models at 1760 km.
For the R1 wave in Fig. 9 (b) and (e), there are low magnitude positive covariances for all length scales, except at 1760 km where the RAL model indicates large positive covariance, and GAL a small negative covariance value. This occurs because at 1760 km, this is almost exactly at the scale of the composite R1’s wavelength of around 2000 km. In RAL, the interscale energy transfer at this length scale is strong and in phase with the wave. The reason this doesn’t coincide with high correlation values (around 0.3 at this length scale) is that the composite comprises a vertical dipole in one phase of the wave, meaning that although upscale energy transfer dominates the composite, the small area of anti-correlated downscale in the wave peak at around 300 hPa reduces the correlation. Correlations for just the upper troposphere-lower stratosphere region are on the order of 0.9 at this length scale in the R1 wave.
@@
For the GAL however, this region of upscale is not nearly as strong, and is out of phase with the wave (hence why the correlation is now negative too). Correlations at length scales 220 km and below are very low and positive in the RAL and low and either side of 0 in the GAL. Correlations between the R1 composite vorticity structure and the interscale energy transfer are high for length scales between 440 and 880 km. Correlations are higher in RAL for the larger length scales. Both models exhibit a correlation close to zero at the largest length scale of 3410 km. This is larger than the wavelength of the wave composite.
@@
In the WMRG wave in Fig. 9 (c) and (f), covariances between WMRG meridional wind and interscale energy transfer are consistently positive for all length scales, and higher in RAL compared to GAL. The wavelength of the WMRG wave in this study period is large (around 7000 km), so the maximum interscale energy transfer length scale considered is at $\lambda/2$. The relationship is the same for the correlations, with a positive trend as the length scale of interscale energy transfer increases towards the scale of the wave. Analysis on a larger domain would be required to extend this analysis beyond 3410 km, since the northern boundary of the CTC is approximately 3000 km from the equator.
@@
For the Kelvin and R1 waves, where the length scales of interscale energy transfer considered begin to approach the scale (wavelength) of the wave composites, the correlations are largest in the range $\lambda/8,\lambda/2]$. At these particular length scales, upscale transfer (deep convection) is often in phase with the dynamical wave. Convection coupled to equatorial waves exhibiting length scales in this range has been qualitatively observed (e.g. Yang et al. 2007a), but here we are able to show quantitatively that the strongest convective coupling occurs across the 440 - 1760 km range of scales. One interpretation of this that requires further investigation is that energy is transferred straight from the mesoscale of 440 km to the synoptic scale of 1760 km (at about $\lambda/2$).
@@
The other interesting comparison we can make here is that this relationship between convection on scales $\lambda/8$ to $\lambda/2$ and the dynamical wave at scale $\lambda$ is redolent of a fundamental property of inverse cascades in 2D turbulence. The idealised study of (Chen et al., 2006) that states that “…energy flux across a length $L$ in the inverse [upscale] cascade range is due to thinning of subscale vortices 4-8 times smaller than $L$ by strain at length scale LL…”. Storer et al. (2023) also found that energy transferred upscale across a mesoscale in the ocean was principally deposited at a scale 3.6 times larger. To our knowledge, this study represents the first instance of a similar result in the atmosphere. This relationship occurs more coherently in RAL than in GAL, with higher correlations and higher covariances between the interscale energy transfer and the wave at these length scales.

### v2

@@ 4.1 Simulated waves and precipitation
From this first analysis of the waves and precipitation in the RAL3 and GAL9 models, we have observed that both models simulate more rain than the observations, the RAL3 particularly so. Excessive tropical rain is commonly observed in CPMs (e.g. Howard et al. 2024). The GAL9 model does not capture the high variability of precipitation seen in the tropics, whereas the RAL3 does. Again, this has been previously observed by other CPM studies (e.g. Judt and Rios-Berrios 2021). Regarding equatorial wave propagation, the waves appear comparable on a Hovmöller plot but a closer investigation reveals there are differences in simulated propagation speed between the analysis and models, with both models simulating a too fast Kelvin and WMRG wave, and GAL9 with a higher variability of simulated wave speeds. The Rossby wave speeds in GAL9 are comparable to the analysis, and faster in RAL3 with higher variability. The distribution of wave variability in the study period across the equatorial band and full depth of the troposphere is comparable between all models and the analysis.
@@ 4.2 Simulated upscale energy transfer and dynamics
To get a first indication of the reasons for the differences in simulated propagating convection between the models, time mean maps of interscale energy transfer for $L$ = 440 km at 200 hPa (the top of the tropical troposphere) are shown in Fig. 5 for GAL9 and RAL. This is the level where horizontal divergence is at a maximum, and so we expect to see high values of upscale energy transfer, because of the mechanism explained at the end of this subsection. For the interscale energy transfer analysed in this study, we do not compare the models to the analysis product. This is because it is at a much coarser resolution than the models and is driven by a data assimilation (DA) scheme with parametrized convection. This means we would expect to see low magnitudes of interscale energy transfer in the anlaysis that we would not trust at the lower length scales analysed since this would be too close to the grid scale of the analysis. Furthermore, the presence of DA in the analysis product would not allow for the investigation of propagating features developing on their own for a longer time period without the impact of the DA scheme.
In Fig. 5, the blue colours indicate the energy is being transferred upscale from scales below 440 km to scales larger than this. The red colours indicate that energy is being transferred downscale from scales larger than 440 km to scales smaller than this. One clear and obvious difference between the two is that in RAL3, there is orders of magnitude more upscale energy transfer, and this is reflected by the deep-tropical mean at 200 hPa being net upscale. The mean was taken over 15S to 15N to avoid impact of sub-tropical systems, such as the subtropical jet. The domain 20S-20N is shown to allow for visualisation of the full extent of the ITCZ in Boreal Summer. The values shown are broadly consistent with those computed by Faranda et al. (2018), the only other paper that has studied local interscale energy transfers in the atmosphere. The magnitudes reported here are on the same order as atmospheric turbulent dissipation rates calculated from observations (Nowak et al., 2025).
@@
The ITCZ is visible in Fig. 5 as the blue band of upscale centred around 10N. Similarly, the SPCZ is visible extending west from New Guinea to the other side of the dateline. Both convergence zones are associated with much stronger upscale energy transfer in the RAL3 model. Another key feature in Fig. 5 is the abundant upscale energy transfer in the Indian Ocean. This period coincides with a strong negative IOD event (cf. Table 1), which is associated with enhanced convection in the southeast Indian Ocean (Saji et al., 1999) – a region which is convectively active even without the negative IOD. Once more, this region of upscale energy transfer is orders of magnitude greater in RAL3 than GAL, saturating the colour scale. Another upscale feature which is much larger and more extensive in RAL3 is the deep convection over the Amazon. Panel (c) in Fig. 5 provides further evidence that the upscale energy transfer at 200hPa is located at the tropical convergence zones. These zones are characterised by consistently high precipitation, and are reflected in the map of mean rainfall over the study period in this panel. The precipitation fields are from GPM-IMERG, and we can see from Fig. 5 that the RAL3 model is simulating a strong ITCZ in terms of 200hPa upscale energy transfer.
@@
Fig. 6 shows longitudinally-averaged, time-mean interscale energy transfer (shading, zero contour thick grey line), zonal mean zonal winds (black contour lines), and the mean meridional circulation (vectors). The main features of Fig. 6 which connect back to Fig. 5 are the sections of upscale energy transfer at the top of the troposphere centred around 200 hPa. This area of upscale is associated with time-average flow and divergence in the tropical upper troposphere as a result of deep convection and associated vertical velocities in the ITCZ centred around 10N. Near the surface, downscale transfer is seen associated with convergence as part of the ITCZ circulation. The RAL3 exhibits both stronger upscale and downscale energy transfer associated with the ascending branch of this circulation, suggesting that explicit representation of convection leads to RAL3 simulating interscale energy transfer that is significantly larger than that in the GAL9 model. The large scale Hadley circulation is of similar strength in both RAL3 and GAL, but RAL3 shows an enhanced secondary circulation at 500 hPa associated with convection. Centred around 5S and 150 hPa is a region of upscale energy transfer which can be identified with the region of strong upscale energy transfer over the Indian Ocean seen in Fig. 5. This region of strong upscale is also associated with downscale energy transfer at low levels underneath the deep convection associated with net flow convergence. As also seen in Fig. 5, this region is about twice as large in RAL3 compared to GAL. This is further evidence of the CPM’s ability to simulate strong upscale energy transfer associated with deep convection. At the southern edge of (a) in the upper troposphere, the signal of the subtropical jet is seen associated with strong zonal winds and strong downscale energy transfer.
@@
In the middle troposphere, there are two areas of weak upscale energy transfer in the RAL3, but only one in the GAL. After consulting maps of time-mean interscale energy transfer at 500 and 600 hPa (not shown), the feature at this level in the RAL3 around 5-10N is associated with the stronger secondary ITCZ circulation in this model, and the presence of cumulus congestus associated with the West African Monsoon. This upscale feature is not present in the GAL. At around 700 hPa between 5 and 10S is another upscale feature which appears in both models. On consultation of the maps at this level this appears to be collocated with upscale energy transfer over the Amazon and Congo rainforests, possibly associated with shallow convection in these regions.
@@
Moving away from the time-mean picture, a Hovmöller of interscale energy transfer is shown in Fig. 7. The contour limits are 5 times higher for the 200 hPa panels than the 850 hPa panels due to the larger overall amount of kinetic energy due to higher wind speeds at this altitude. For the region shown in Fig. 7 spanning the Indian Ocean, Maritime Continent, and West Pacific, there are abundant eastward-propagating downscale energy transfer features at 850 hPa which are broadly collocated with the propagating rainfall and Kelvin waves in Fig. 3. These downscale features are occasionally coincident with upscale energy transfer at 200 hPa (e.g. see in the RAL3 starting at 90E on 2016-08-01), indicative of propagating deep convection (likely a CCKW). The picture at 200 hPa is mostly dominated by westward-propagating features. These Hovmöller diagrams were analysed with waves overlain and for symmetric, anti-symmetric, and off-equatorial interscale energy transfer (not shown). We found that the westward-propagating features were indeed indicative of CCEW activity: R1 waves in the case of the off-equatorial interscale energy transfer (0-20N average), and WMRG waves for the anti-symmetric part (calculated by taking the average of the difference between the fields in the 20 degrees north and south of the equator).
Figure 7: Time-longitude (Hovmöller) diagrams of $\mathcal{D}_{L}(\mathbf{u})$ for $L$ = 440 km averaged between 10S and 10N for the 20 day study period, limited to 45 - 180E to remove the strong orographic effects of the East African Highlands at 850 hPa. (a) for GAL9 at 200hPa; (b) RAL3 at 200 hPa; (c) GAL9 at 850 hPa, and (d) RAL3 at 850 hPa. Note the contour limits are 5 times higher in (a) and (b) than in (c) and (d). Colour-coded numbers in the top right of each panel are mean values of the data shown with the colour indicating net upscale (blue) or net downscale (red). The dashed line indicates the time of the snapshot profile shown in Fig. 8. The time mean has been removed to clarify the propagating features, and this has been done for Fig. 8, and for calculating the composites shown in Fig. 9.
@@ 4.3 Equatorial waves and upscale energy transfer
The picture of the imprint of the general circulation on the interscale energy transfer can be seen in Fig. 7 by the fact that the 200 hPa regions are net upscale and the 850 hPa regions net downscale. Once again it is evident that RAL3 is simulating significantly more upscale than GAL9 at 200 hPa – 2 orders of magnitude more in this case. To investigate how propagating features in Fig. 7 are associated with different equatorial waves, longitude-pressure snapshots at the time indicated by the dashed line in Fig. 7 of dynamical wave fields and interscale energy transfer are shown in Fig. 8. All wave snapshots are shown across the longitudes where they are strongest in this season. For the KW this is over the Indian Ocean, for the R1 wave it is over the Atlantic, and for the WMRG over the West Pacific.
From figures 8 (a) and (d), we can see upscale energy transfer related to Kelvin wave divergence. Upscale energy transfer in the upper troposphere in the divergent phase of the wave is observed because the convecting plume coupled to the wave is expanding near the tropopause, transferring energy from scales smaller than 440 km (the scale of the updraft) to scales larger than 440 km (toward the scale of the wave). This is balanced by downscale energy transfer in the convergent phase of the wave, both beneath the divergence in the lower troposphere, and laterally in the wave troughs either side of the upper-level wave divergence. Another thing to note from Fig. 8 is that the vertical structure of the Kelvin wave in RAL3 is closer to observations and theory than GAL9 (cf. Straub and Kiladis 2003; Yang et al. 2007a). The wave is higher amplitude, exhibits clear westward tilt with height, and upward propagation into the stratosphere in RAL3, whereas the GAL9 does not exhibit these qualities as clearly, especially in the lower troposphere.
@@
The composites in Fig. 9 provide more information than the instantaneous structures seen in Fig. 8, demonstrating the systematic features that are present over time in interscale energy transfer and dynamical equatorial wave fields. Regarding the Kelvin wave and symmetric eastward-propagating interscale energy transfer, it can be seen that the dynamical wave composite is similar in both the RAL3 and the GAL9 models. As expected, there is more extreme magnitudes of upscale energy transfer in the RAL3, and this appears to be more highly correlated with the KW divergence in the upper troposphere than the GAL. The correlations between the propagating interscale energy transfer and corresponding equatorial waves for varying $L$ are outlined in Fig. 10.
@@
Figures 9 (b) and (e) show again the positive correlation between R1 vorticity and interscale energy transfer. This positive correlation between off-equatorial interscale energy transfer and R1 vorticity is reflected in Fig. 10, with correlations in RAL3 being on average 10% higher than in GAL, and more consistent (same sign for all $L$).
@@
The WMRG composites span a greater longitudinal extent than the other waves in Fig. 9 (c) and (f). This is to allow room to show the striking differences in their simulation between the models. In the RAL3 there is the characteristic vertical propagation of the WMRG wave typically seen during the westerly phase of the QBO (Yang et al., 2012). The QBO is indeed in a westerly phase during DYAMOND Summer, see Table 1. No such vertical propagation is seen in the GAL. The GAL9 does however have a slightly stronger baroclinic wave structure, even if the wave is much weaker. The composite is again consistent with the snapshot for RAL3, with southward WMRG winds being collocated with anti-symmetric upscale energy transfer. This collocation is also present in GAL9 albeit with lower values of interscale energy transfer.
@@
Covariance and correlations between symmetric eastward, off-equatorial westward, and anti-symmetric westward interscale energy transfer and the dynamical waves are presented for varying length scales in Fig. 10. The correlations and covariances displayed here are comparing the dynamical wave and interscale energy transfer composites shown in Fig. 9 for multiple length scales. For reference, the fifth data points for each line represent the $L=440$ km length scale plotted in Fig. 9.
@@
Figure 10: Covariances and correlations between the fields shown in Fig. 9 as a function of characteristic length scale $L$ . Kelvin wave divergence about the equator is compared with equatorially symmetric KW-following interscale energy transfer, R1 off-equatorial vorticity with off-equatorial R1-following interscale energy transfer, and WMRG cross-equatorial $\upsilon$ with anti-symmetric WMRG-following interscale energy transfer, as in Fig. 9. Correlations and covariances for GAL9 in cyan, for RAL3 in orange. The blue vertical lines represent multiples of the composite dominant wavelength, with the dotted line $\lambda/8$, dot-dash $\lambda/4$, dashed $\lambda/2$, and solid line at the dominant wavelength, $\lambda$.
@@
For the Kelvin wave in Fig. 10 (a) and (d), we can see that the covariance between the dynamical wave and interscale energy transfer is stronger in the RAL3 for length scales 880 km and lower, comparable and at a maximum for $L=1760$ km, and then the other side of the zero line for the maximum length scale analysed here of 3410 km. Consultation of the composites at these two large length scales (not shown) reveals that the reason for this is because at $L=1760$ km, the regions of strong interscale energy transfer are close to the scale of the wave (between $\lambda/4$ and $\lambda/2$) and almost in-phase with them. This means a correlation of over 0.5 in both models, and with a highly negative covariance - negative (upscale) collocated with positive (divergent) wave-dynamical flow. However at the 3410 km length scale, the interscale energy transfer features are double the size, and are almost large enough to stretch across an entire wavelength with the same sign, leading to correlations close to 0 and a positive covariance, since the deepest regions of upscale energy transfer are located in the convergent phase of the wave. Correlations are strongest for $L=880$ km. The reasons for this will be discussed at the end of this section in conjunction with the information from the other waves. Correlations are slightly larger in magnitude in the RAL3 for 440 and 880km, and comparable for both models at 1760 km.
@@
For the R1 wave in Fig. 10 (b) and (e), there are positive covariances for all length scales, except at 1760 km where the GAL9 model negative covariance. This occurs because a length scale of 1760 km is almost exactly at the scale of the composite R1’s wavelength (around 2000 km). In RAL3, the interscale energy transfer at this length scale is large and in phase with the wave, whereas in GAL9 it is out of phase, hence the negative covariance and correlation. Correlations at length scales of 220 km and below are small and positive in both models. Correlations between the R1 composite vorticity structure and the interscale energy transfer are highest for $L=880$ km in RAL3, and at $L=440$ km in GAL9. This represents close to a half and a quarter of the wavelength, respectively.
@@
In the WMRG wave in Fig. 10 (c) and (f), covariances between WMRG meridional wind and interscale energy transfer are consistently positive for all length scales, and higher in RAL3 compared to GAL. The wavelength of the WMRG wave in this study period is large (around 7000 km), so the maximum interscale energy transfer length scale considered is at $\lambda/2$. The relationship is the same for the correlations, with a positive trend as the length scale of interscale energy transfer increases towards the scale of the wave. Analysing larger length scales would not make sense for this study since the kernel is on a spherical cap, and for large synoptic or planetary scales, information from subtropical and even midlatitude motions would be included, meaning the interscale energy transfer values would not be meaningfully connected to the equatorial wave of interest. This is acknowledged as a limitation of this study, but could be ameliorated if a search was conducted for lower wavelength WMRG waves in future work.
@@
For the Kelvin and R1 waves, where the length scales of interscale energy transfer considered begin to approach the scale (wavelength) of the wave composites, the correlations are largest in the range $\lambda/8,\lambda/4]$. At these particular length scales, upscale transfer (deep convection) is often in phase with the dynamical wave. Convection coupled to equatorial waves exhibiting length scales in this range has been qualitatively observed (e.g. Yang et al. 2007a), but here we are able to show quantitatively that the strongest convective coupling occurs close to the 880 km length scale. One interpretation of this that requires further investigation is that energy is transferred straight from the large mesoscale of organised convection at 880 km to the wave scale at about 4000 km.
@@
The other interesting comparison we can make here is that this relationship between convection on scales $\lambda/8$ to $\lambda/4$ and the dynamical wave at scale $\lambda$ is redolent of a fundamental property of inverse cascades in 2D turbulence. The idealised study of (Chen et al., 2006) that states that “…energy flux across a length $L$ in the inverse [upscale] cascade range is due to thinning of subscale vortices 4-8 times smaller than $L$ by strain at length scale LL…”. Storer et al. (2023) also found that energy transferred upscale across a mesoscale in the ocean was principally deposited at a scale 3.6 times larger. To our knowledge, this study represents the first instance of a similar result in the atmosphere. This relationship occurs more coherently in RAL3 than in GAL, with higher correlations and higher covariances between the interscale energy transfer and the wave at these length scales.